OpenStudy (anonymous):

simplfy rational algebraic expression x-5/(5x-sqrt5)(5x+sqrt5)

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

is this \[\frac{x-5}{(5x-\sqrt{5})(5x+\sqrt{5})}\]

OpenStudy (anonymous):

ok here is what you use: \[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

therefore you denominator is straightforward. it is \[(5x)^2-\sqrt{5}^2\] aka \[25x^2-5\]

OpenStudy (anonymous):

yep..then..

OpenStudy (anonymous):

stick that under your numerator and be done!

OpenStudy (anonymous):

ok il try it..thanks...